OpenStudy (anonymous):

Find the distance between the points (2, -5) and (-3, 0).

OpenStudy (anonymous):



OpenStudy (anonymous):

what are those two distances?

OpenStudy (anonymous):

isn't it just a formula

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

OpenStudy (anonymous):

2-3 5-0

hartnn (hartnn):

it will be 2-(-3) = 2+3 = ? and -5-0 = ?

OpenStudy (shaik0124):

d=root50

OpenStudy (anonymous):